OpenStudy (highschoolmom2010):

A parallelogram has sides measuring 23.8 m and 35.3 m. The height corresponding to the 23.8-m base is 9.9 m. Find the height, to the nearest tenth of a meter, corresponding to the 35.3-m base

OpenStudy (rational):

hint : area

OpenStudy (nurali):

he area of parallelogram is : side times height corresponding to that side, so 23,8*9,9=35,3*h h=6.6764

OpenStudy (austinl):

 Is this the figure?

OpenStudy (rational):

spoiler would be good if we let the asker do the problem on her own :o

OpenStudy (rational):

thats for nurali

OpenStudy (highschoolmom2010):

235.62=35.3h it didnt have a figure

OpenStudy (highschoolmom2010):

h=6.67

OpenStudy (austinl):

You would find the area of the figure, which is the base times the height. You can do so by using the base and height given. Then you need to use that area and set it equal to the second base and an arbitrary constant for the second height, which you have done. Then you solve for the arbitrary constant. Which you also just did correctly. \(h_{(35.3)}=6.67479\)

OpenStudy (rational):

remember you need to round it to nearest tenth not hundredth

OpenStudy (highschoolmom2010):

h=6.7

OpenStudy (austinl):

Didn't notice that part of the instructions, good catch @rational And yes, that is correct.

OpenStudy (rational):

\(\large \color{red}{\checkmark}\)

OpenStudy (rational):

oops ! i meant for highschoolmom austin :)